OpenStudy (anonymous):

The distance, d, an object travels is represented by the formula d = rt, where r is the rate at which the object is traveling and t is the duration of time over which the object travels. If d has units of meters and r has units of meters per minute, what must be the units of t? A. meters B. seconds C. minutes D. meters per second

OpenStudy (anonymous):

@myko @Isaiah.Feynman @e.mccormick @mathgeek27

OpenStudy (anonymous):

@mathmath333

OpenStudy (isaiah.feynman):

This should be intuitive.

OpenStudy (isaiah.feynman):

What is time measured in?

OpenStudy (mathmath333):

\(\large\tt \color{black}{r=\dfrac{d}{t}}\) \(\large\tt \color{black}{(unit)t=\dfrac{meter}{\dfrac{meter}{minute}}}\) \(\large\tt \color{black}{(unit)t=minute}\)